OpenStudy (anonymous):

Please help, I will give best response. It's a practice test, so it's not cheating if you help. create 2 union sets that have the union of {-1,0,1,2,3,4,5,6}

Directrix (directrix):

This question was asked earlier today. There are lots of different answers that are correct. The union is the combination of two sets. http://openstudy.com/users/directrix#/updates/56122744e4b0023ff68954c6

Directrix (directrix):

This person chose" >>is this a good one? A= -1,0,1,2,3 B=3,4,5,6 Yes, 3 appears in both sets but in the union, only one 3 would be included. Write it as the following: {-1,0,1,2,3} U {3,4,5,6} --> Choose something different. Just be sure all the numbers are in one or both of the two sets which will be "unionized." Post what you get and I'll check it for you if you want.

MsBrains (ms-brains):

I just want somebody to explain how I am supposed to figure this out.

Directrix (directrix):

You first have to know what the union of two set is, the definition. The union of two sets A and B is the set of elements which are in A, in B, or in both A and B. In symbols, . For example, if A = {1, 3, 5, 7} and B = {1, 2, 4, 6} then A ∪ B = {1, 2, 3, 4, 5, 6, 7}.

Directrix (directrix):

In your problem, you are asked to work backwards. That is, you take the answer (the union set) and give two sets which could have the given numbers in the union of the two sets. There are many different answers for A and B that you could give. When you take the union, you have to get what is given as the union set in the problem.

MsBrains (ms-brains):

Thanks so much @Directrix can you help with :Create 2 sets that have and intersection of {2,4,6}

Directrix (directrix):

In mathematics, the intersection A ∩ B of two sets A and B is the set that contains all elements of A that also belong to B (or equivalently, all elements of B that also belong to A), but no other elements.

Directrix (directrix):

Make up two sets, A and B, such that each set contains the elements 2, 4, and 6. The sets may contain other numbers but each of A and B sets much have 2, 4, and 6, among the elements. There are many different answers. Such as, A = { 1, 2, 3, 4, 5, 6} and B = { -3, 0, 2, 4, 5, 6, 10}

MsBrains (ms-brains):

3.Create 2 sets whose INTERSECTION results in a null set. Last question, I have to prepare for the test.

Directrix (directrix):

A = { 1, 2, 3, 4, 6} n { -3, 0, 2, 4, 5, 6, 10} = {2,4,6} Bad answer alert. I had to take the 5 out of set A because set B had a 5 in it and the intersection would have contained a 5 which is not allowed in the problem.

Directrix (directrix):

The null set is the empty set. Two sets with no common elements would have an intersection that is the null set. For example, A = { 1, 2} and B = { 7, 8} A n B = { } which is how I write the empty set because { } is a set with no element.

MsBrains (ms-brains):

Honestly though, thank you so much man! :) You're a life saver.

Directrix (directrix):

You are welcome.

MsBrains (ms-brains):

You don't by any chance know how simplify expressions, do you? That's the next in my Algebra study book..

MsBrains (ms-brains):

Simplify both expressions 7+2(4x-2)+10 @Directrix

MsBrains (ms-brains):

And please show me how you solved it

Directrix (directrix):

7 + 2(4x-2) + 10 = 7 + 2*4x -2*2 + 10 = 7 + 8x - 4 + 10 = 8x + 13

MsBrains (ms-brains):

Simplify 10+2+11xy squared2 +3y Squared2 +3y Squared3 +7x _6xy Squared2 -4x-2y sqaured3? It's kinda tricky, and if you don't know to solve it's okay.. :\

MsBrains (ms-brains):

@Directrix